OpenStudy (anonymous):

Where does the graph of y = 2x2 - x - 3 cross the x-axis?

sam (.sam.):

You know that when the graph touches the x-axis the 'y' values is zero

sam (.sam.):

So, 2x^2 - x - 3=0 (x+1)(2x-3)=0 Set each brackets =0 and solve for x